Breakin Away Company has three employees-a consultant, a computer programmer, and an administrator. The following payroll information is available for each employee:

Consultant Computer Programmer Administrator
Regular earnings rate $2,110 per week $28 per hour $50 per hour
Overtime earnings rate Not applicable 1.5 times hourly rate 2 times hourly rate
Number of withholding allowances 3 2 1

For the current pay period, the computer programmer worked 60 hours and the administrator worked 50 hours. The federal income tax withheld for all three employees, who are single, can be determined by adding $356.90 to 28% of the difference between the employee's amount subject to withholding and $1,796.00. Assume further that the social security tax rate was 6%, the Medicare tax rate was 1.5%, and one withholding allowance is $70.

Determine the gross pay and the net pay for each of the three employees for the current pay period. Assume the normal working hours in a week are 40 hours. If required, round your answers to two decimal places.

Consultant Computer Programmer Administrator
Gross pay $ $ $
Net pay $ $ $
Answer #1

Calculate gross pay as follows:

A B C D E F G
1 Total hours worked Regular hours Overtime hours Regular pay Overtime pay Gross pay
2 Consultant (per week) 2110
3 Computer Programmer 60 40 20 1120 840 1960
4 Administrator 50 40 10 2000 1000 3000

Above figures are calculated in the following manner:

Total hours worked Regular hours Overtime hours Regular pay Overtime pay Gross pay
Consultant (per week) 2110
Computer Programmer 60 40 20 =C3*28 =D3*28*1.5 =E3+F3
Administrator 50 40 10 =C4*50 =D4*50*2 =E4+F4

Calculate net pay as follows:

A B C D E F G H I J
1 Gross pay One withholding allowance Number of withholding allowances Total withholding allowance Amount subject to withholding Federal Income Tax Withheld Social Security Tax Withheld Medicare Tax Withheld Net Pay
2 Consultant (per week) 2110 70 3 210 1900 386.02 126.60 31.65 1565.73
3 Computer Programmer 1960 70 2 140 1820 363.62 117.60 29.40 1449.38
4 Administrator 3000 70 1 70 2930 674.42 180.00 45.00 2100.58

Above figures have been calculated in the following manner:

Gross pay One withholding allowance Number of withholding allowances Total withholding allowance Amount subject to withholding Federal Income Tax Withheld Social Security Tax Withheld Medicare Tax Withheld Net Pay
Consultant (per week) 2110 70 3 =C2*D2 =B2-E2 =356.9+((F2-1796)*28%) =B2*6% =B2*1.5% =B2-G2-H2-I2
Computer Programmer 1960 70 2 =C3*D3 =B3-E3 =356.9+((F3-1796)*28%) =B3*6% =B3*1.5% =B3-G3-H3-I3
Administrator 3000 70 1 =C4*D4 =B4-E4 =356.9+((F4-1796)*28%) =B4*6% =B4*1.5% =B4-G4-H4-I4